### v2

# Large Multimodal Models for Embodied Intelligent Driving: The Next Frontier in Self-Driving?

Abstract
The advent of Large Multimodal Models (LMMs) offers a promising technology to tackle the limitations of modular design in autonomous driving, which often falters in open-world scenarios requiring sustained environmental understanding and logical reasoning. Besides, embodied artificial intelligence facilitates policy optimization through closed-loop interactions to achieve the continuous learning capability, thereby advancing autonomous driving toward embodied intelligent (El) driving. However, such capability will be constrained by relying solely on LMMs to enhance EI driving without joint decision-making. This article introduces a novel semantics and policy dual-driven hybrid decision framework to tackle this challenge, ensuring continuous learning and joint decision. The framework merges LMMs for semantic understanding and cognitive representation, and deep reinforcement learning (DRL) for real-time policy optimization. We starts by introducing the foundational principles of EI driving and LMMs. Moreover, we examine the emerging opportunities this framework enables, encompassing potential benefits and representative use cases. A case study is conducted experimentally to validate the performance superiority of our framework in completing lane-change planning task. Finally, several future research directions to empower EI driving are identified to guide subsequent work.
I Introduction
Autonomous driving represents the new generation of intelligent transportation that can improve safety and operational efficiency significantly [1]. The core of this change is that self-driving vehicles accurately perceive their dynamic environment and take driving decisions in real time. However, due to the lack of sustained environmental understanding and the absence of logical reasoning capability [2], the limitations of conventional modular design approaches are becoming more apparent. Under these shortcomings, the reliability, safety, and generalization performance of autonomous driving systems are critically constrained in real-world complex traffic scenarios.
The field of artificial intelligence (AI) has been revolutionized by Large Language Models (LLMs) and the recently emerged Large Multimodal Models (LMMs) [3]. In response to the above challenges, LMMs have become an essential technology to improve the environmental understanding and decision-making ability of autonomous driving systems. By interpreting multimodal driving data via LMMs, self-driving vehicles acquire a thorough semantic understanding of complex scenes, and perform behavioral prediction [4], thereby enhancing their safety and reliability in real-world operations. Several recent works on LMM-empowered autonomous driving have been explored from the different perspectives, such as external environment understanding [5], driving decisions and control [6], driver state recognition [7], etc. However, as the traffic environment becomes increasingly complex, these capabilities presented by LMMs struggle to support autonomous adaptation and evolution in real-time interactions between self-driving vehicles and their environment under dynamic conditions. Therefore, more powerful methods are needed to address the growing complexity of traffic environments.
In this context, embodied AI has emerged as a revolutionary means of promoting autonomy in autonomous driving systems. Unlike conventional AI models, embodied AI emphasizes the interactions between embodied agents and their physical environment [8]. By embedding embodied AI, self-driving vehicle can adapt to dynamic environmental changes and make intelligent decisions based on the interactive data. Such adaptability enables the vehicles to continuously optimize and evolve their driving policies through interaction in complex environment, thereby realizing the embodied intelligent (EI) driving paradigm. Currently, embodied AI is in the exploratory stage, used for real-time decision-making and action planning in dynamic and uncertain environment. Great potential has been particularly demonstrated in the transportation-related scenarios, including vehicular networks and rail transportation systems [9, 10, 11, 12].
TABLE I: Comparison with the existing works on LMMs and DRL in the autonomous driving systems and other transportation-related scenarios.
Reference Year LMMs DRL Main Contributions
Liao et al.[5] 20252025 $\checkmark$ $\times$ They proposed an LMM-based traffic accident anticipation framework through the CoT prompting. By hierarchically fusing video, optical flow, and linguistic descriptions, the early perception capability can be effectively enhanced for high-risk scenarios.
Xu et al.[6] 20242024 $\checkmark$ $\times$ They proposed an interpretable end-to-end autonomous driving system based on the visually instruction-tuned and mixed fine-tuned LMMs. This system achieves the unified modeling from forward-view videos to low-level control commands and natural language explanations.
Hu et al.[7] 20252025 $\checkmark$ $\times$ They proposed a training-free multimodal LLM method driven by human-centric context and self-uncertainty. This method extracts the scene graphs and integrates the multiple reasoning responses through the evidence theory to recognize the driver states accurately without any training.
Zhang et al.[9] 20252025 $\checkmark$ $\checkmark$ They proposed an embodied AI-enhanced vehicular network framework that integrates the vision-LLMs for semantic extraction and the DRL to stabilize decision-making under uncertainty. This framework significantly enhances the efficiency of semantic communication via the vision-LLMs and decision‑making through the DRL.
Chen et al.[10] 20252025 $\checkmark$ $\times$ They developed an embodied AI‑enabled perception and decision-making framework for the scenarios of connected vehicles. By employing LMMs, this framework allows vehicle-to-vehicle (V2V) data to achieve greater situational awareness and reliable decision-making in uncertain scenarios.
Zhou et al.[11] 20242024 $\checkmark$ $\checkmark$ They proposed a framework for an autonomous operation of rail transportation system enabled by embodied AI through integrated perception, execution, and learning. This framework allows the system to improve its adaptability to the environment and enhance autonomous decision-making capabilities.
Li et al.[12] 20242024 $\checkmark$ $\times$ They introduced an LMM-empowered embodied AI framework for autonomous driving in mining. By incorporating the interactive learning and CoT reasoning, the framework can improve the decision interpretability and environmental adaptability of mining vehicles.
This article $\checkmark$ $\checkmark$ We propose a joint decision framework by merging the LMMs and DRL for EI driving. Our framework particularly combines the multimodal semantic reasoning with closed-loop policy optimization. Such combination enhances adaptation to dynamic driving environments and the overall generalization.
Although these achievements are of great significance, as shown in Table I, a closed-loop design mechanism capable of continuous learning in response to environmental changes within dynamic traffic conditions has yet to be established. While LMMs enhance understanding of multimodal driving data, deep reinforcement learning (DRL) is required to facilitate joint decision-making in dynamic environment. The decision process often necessitates continuous interaction between EI vehicle and its environment. Therefore, incorporating the continuous learning and joint decision in dynamic scenarios into EI driving remains an emergent research frontier.
In this article, we propose employing the joint decision based on LMMs and DRL to enhance EI driving. LMMs inherit and extend the capabilities of LLMs, traditionally designed for textual data processing. Through architectural adaptation and multimodal collaborative training via LMMs, EI vehicle effectively handles the collected multimodal data [13]. These data processed and integrated into internal representation thereby enhance its generalization ability to new environment and tasks. In contrast, DRL behaviors can be adjusted by real-time feedback, which results in adaptive policy development in dynamic operations [14, 15]. In DRL, the agent interacts with the environment and as a result, receives feedback in the form of rewards or penalties. And thus, the agent learns how to make optimal decisions. We will now demonstrate the application of LMMs and DRL-based joint decision across various scenarios, as shown in Fig. 1. The contributions of this article are summarized below.
Fundamental Concepts and Requirements. We introduce the foundational concept of EI driving, highlighting its closed-loop pipeline of perception, decision, and execution as a core architectural design for EI vehicle. A brief outline of the basic architecture of LMMs, including modules and implementation process, is provided. Given such fundamentals, we demonstrate the requirements and necessities for integrating LMMs with EI driving.
Framework and Opportunities. To implement the LMM-enhanced EI driving, we propose a semantics and policy dual-driven hybrid decision framework, which integrates LMMs, for semantic understanding and cognitive representation, and DRL, for real-time policy optimization. Additionally, we identify the emerging opportunities this framework enables, including the potential benefits and transformative use cases.
Case Study and Evaluations. We evaluate the proposed framework in a case study of lane-change planning, utilizing the scenario mixed with an EI vehicle and multiple conventional human-driven vehicles. The experimental results indicate that our framework achieves superior performance in policy quality and dynamic adaptability.
Figure 1: An illustration of the EI driving scenarios enabled by the joint decision approach based on LMMs and DRL. Such approach shows great promise in three illustrative application scenarios: i) Urban robotaxi, which is on-demand and driverless transport, faces tremendous challenges in negotiating highly dynamic traffic scenarios. Examples would be predicting a pedestrians intent, and avoiding a sudden obstacle in a narrow corridor. ii) Autonomous freight operates in trunk-line logistics, ports, and industrial parks, where good performance is needed in complex geometrical road situations. Representative scenarios include container handling, and narrow space delivery robots. iii) Structured shuttle, driving in structured and semi-open environments, must achieve precision docking. Typical scenarios include airport shuttle service, and micro-circulation bus.
II An Overview of LMMs and EI Driving
This section begins with the fundamental concept of EI driving, followed by the general framework of LMMs. Then, the necessities of integrating EI driving and LMMs is discussed.
II-A Fundamentals of EI Driving
Embodied AI refers to the intelligent behavior generated by the interaction between embodied agent and its external environment [8]. Such behavior emphasizes the direct physical world interaction, and the close coupling among perception, action, and learning. By integrating embodied AI, embodied agent obtains human-like environmental perception and real-time feedback. This enables the embodied agent to demonstrate the capabilities of continuous learning and autonomous evolution through its morphological characteristics and dynamic interaction with the environment.
Figure 2: The architecture of EI vehicle, which consists of the embodied perception module, embodied decision module, and embodied execution module. The EI vehicle first achieves the autonomous environmental perception through multimodal sensors, and collects heterogeneous data sources, such as vision, radar, velocity, and textual data. Based on the comprehensively fused perception, the EI vehicle proceeds to interpret the surrounding environment and perform the scenario-based decision-making, thereby formulating the high-level policy intentions. Finally, these policy intentions are translated into the concrete control actions for EI vehicle. These modules operate as a closed-loop pipeline, including the perception, decision, and execution.
The integration of embodied AI into self-driving vehicle enhances its ability to comprehensively understand all interactive behaviors, thereby showing a higher degree of intelligence. By continuously collecting multimodal data through environment interaction and feeding the data back to EI vehicle, driving maneuver commands are rapidly generated and executed, enabling real-time adaptation to dynamic environmental conditions. As shown in Fig. 2, the architecture of EI vehicle adopts a closed-loop structure of perception, decision, and execution [11], corresponding to the following three modules.
Embodied Perception. Based on onboard sensors, the embodied perception module collects high-fidelity, multimodal raw data from physical world, endowing EI vehicle with the capability for active understanding of complex environment. According to data sources, the collected data include four modalities: vision, radar, velocity, and text. The visual modality utilizes onboard camera to acquire the environmental images, which facilitates the perception of traffic elements, such as roads, obstacles, and traffic signs. The radar modality uses the LiDAR and radar sensors to achieve precision detection of dynamic objects. The velocity modality integrates the GPS and wheel-speed sensors to obtain real-time position and motion states of EI vehicle. The textual modality analyzes the system internal information and interprets the driver’s instructions.
Embodied Decision. As a modular core of EI vehicle, the embodied decision module efficiently converts perceptual results into the executable behavioral policies. By continuously evaluating real-time traffic conditions, the module dynamically adjusts its policy to jointly optimize the cognition and decision. Through various intelligent decision-making methods, e.g., LMMs and DRL algorithms, the dynamic policy coupled with the vehicle’s physical properties and external environment is generated. Besides, real-time evaluations of task requirements, environmental changes, and vehicular capabilities are performed. This enables the flexible responses to emergencies, and the formulation of safe yet operationally efficient action plans. The mechanism of self-learning and optimization is also integrated into the module, realizing the continuous evolution during actual operation. The adaptability and overall intelligence level are thereby enhanced for decision-making process.
Embodied Execution. The embodied execution module translates executable decisions into precise driving maneuver commands, which embodies precision control and action execution. With the help of a closed-loop coordination with the embodied perception and decision modules, maneuver command dispatching like driving, steering and braking can be accurate. Besides, the module highlights adjusting actions and the external environment with feedback in real-time. The actuation system of EI vehicle should be highly responsive and operationally robust. The feedback data enables dynamic fine-tuning and optimization of the execution policies, thus guaranteeing smooth motion and safe operation.
Figure 3: An overview of the overall framework of LMMs. In particular, the LMMs achieve unified cross-modal understanding and support any-to-any modality inputs and outputs through multimodal encoders, multimodal alignment learning, multimodal generation, and multimodal instruction tuning.
II-B Fundamentals of LMMs
LLMs learn from massive amounts of text to consider global information within input sequences and target the most relevant parts. LMMs utilize LLMs as their cognitive core, by integrating the modality-specific encoders and various decoders [13]. This thereby achieves the heterogeneous multi-source data processing and the generation of multimodal outputs. Notably, LMMs adopt a two-stage training method: multimodal pretraining and fine-tuning. Pretraining stage bridges the semantic gap between multimodal features and textual features to achieve cross-modal alignment. Fine-tuning stage further enhances their capabilities to understand and execute complex instructions. As shown in Fig. 3, the LMM framework includes the following key components.
Multimodal Encoders. Multimodal encoding aims to extract discriminative features from heterogeneous input data. For example, the encoders can extract the visual features from visual modalities (e.g., images or videos), and the acoustic features from auditory modalities (e.g., audio waveforms). These features are converted into a unified representation space to achieve effective interaction between cross-modal information. Due to the differences across various data modalities, the features of each modality are encoded differently in the representation space. Thus, there exists a need for feature alignment learning to address this problem.
Multimodal Alignment Learning. Multimodal alignment learning is important to eliminate cross-modal semantic gap. The multi-head attention mechanism is used to focus on discriminative features in different modal subspaces. At each level of attention, certain modality-specific features are captured, which are then integrated into their own higher representation space, thereby making them cross-modally consistent.
Multimodal Generation. Multimodal generation seeks to effectuate a dynamic composition between input and output modalities. Advanced generative methods are increasingly being used in LMMs, including Transformers and diffusion models. Different input data modalities can give rise to semantic information. The cross modal outputs are then generated in any modality combination.
Multimodal Instruction Tuning. Multimodal instruction tuning aims to comprehend and carry out cross-modal instructions, which refers to building the pairs of instruction semantics to the output data. LMMs applies the In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning capability in a multimodal environment, enhancing their cross-modal generalization performance. In complex and varied cross-modal tasks, performance enhancement is especially observed.
II-C Necessities for Integrating EI Driving and LMMs
EI driving operates in highly complex traffic environment, characterized by heterogeneous multimodal data, multi-scale spatial structures, and dynamic human-machine interactions. The cross-modal understanding and task generalization capabilities of LMMs offer a consolidated groundwork for perception and decision-making at EI vehicle, thereby acting as a vital technological pillar for the realization of EI driving.
Semantic Understanding of Complex Traffic Scenarios. One of the critical challenges in EI driving is how to process heterogeneous multimodal data. Enabled by the cross-modal embedding and semantic unification, LMMs address this challenge by projecting diverse data modalities into a coherent cognitive representation space. This significantly enhances the ability of EI driving system to understand and reason about semantic relationship among environmental elements.
Human-Vehicle Collaboration and Teaming. LMMs can jointly model the natural language, visual behavioral signals, and traffic context. EI driving system thereby infers the intentions of drivers, pedestrians, and passengers through the established semantic relationship. This enables the EI vehicle to interpret and explain its own behavior. With LMMs, the system can respond to the driving maneuver commands and predict the actions of interactive embodied agent, thereby facilitating operationally efficient and interactionally fluent human-machine cooperative driving.
Human-Like Decision-Making in Dynamic Interaction. EI driving system needs to make decision under highly uncertain and rapidly evolving environment. By integrating historical motion trajectories, environmental states, and contextual cues, LMMs perform the execution of CoT reasoning to generate higher-level policies, including both intent understanding and trajectory prediction. The flexible decision-making capability empowers the system to adapt to complex scenarios, thereby advancing its level of generalizable intelligence.
Figure 4: Illustration of the proposed semantics and policy dual-driven hybrid decision framework to implement the LMM-empowered EI driving. Three successive stages are designed and incorporated into this framework: semantic pipeline, policy pipeline, and fusion pipeline.
III Integration of LMMs and EI Driving
This section proposes a procedural framework for implementing the LMM-empowered EI driving. The emerging opportunities this framework enables are also introduced, including the potential benefits and promising use cases.
III-A Framework
To implement the LMM-enabled EI driving, we propose a semantics and policy dual-driven hybrid decision framework, as shown in Fig. 4. The proposed framework integrates DRL, known for its high responsiveness and policy adaptability in complex dynamic environment, with LMMs, which exhibit exceptional capability in semantic reasoning and high-level decision representation. The combination of these complementary strengths empowers EI driving system to understand task context and adapt to dynamic environmental conditions.
Stage 1. Semantic Pipeline. The semantic pipeline leverages LMMs to obtain the higher-level driving action policy, through the CoT processing of unified multimodal cognitive representation. The procedure of this pipeline is specified as follows.
  • Step 1. Feature Encoding: The feature encoding performs the semantic-level encoding operations on raw multimodal data. The inputs from diverse sources, including vision, radar, semantic annotations, and communications, are transformed into the machine interpretable semantic feature representations. This step preserves modality-specific structural information and establishes preliminary cross-modal association. The semantic relationship between different modalities can be thereby learned.
  • Step 2. Data Alignment: Firstly, the cross-modal projection constructs a feature input projection network to achieve consistent mapping of non-linguistic modalities, e.g., images, velocity, and radar signals, into a unified semantic space. This step ensures that all heterogeneous data are coherently represented within a unified cognitive representation space. Then, LLMs comprehensively analyze the driving scenarios by processing text commands and multimodal data. Text outputs and signal tokens are generated for control and decision-making, including path planning and behavior prediction. Finally, the output projection maps signal tokens to specific modal demand space, used for the multimodal content generation.
  • Step 3. Action Realization: The action realization transforms various signal tokens into the decisions and predictions from multimodal data, to guide the action commands of EI vehicle. This step transforms high-level semantic intents into specific physical actions. Visual interface prompts and voice interaction feedback are used to ensure system interpretability and facilitate human-centered interactivity.
  • Step 4. Instruction Tuning: The instruction tuning acquires an understanding of the policy intentions through a limited number of examples. CoT reasoning and ICL mechanisms facilitate hierarchical task decomposition, which progressively derives behavioral logic to enhance the precision of instruction generation. Through continuous reasoning, the EI driving system improves environmental adaptability and task execution stability.
Stage 2. Policy Pipeline. To ensure real-time performance optimization of EI vehicle in dynamic and open environment, the policy pipeline employs DRL approach that directly maps the collected multimodal data into executable action policy. Within this DRL framework, the policy network parameters are continuously optimized through online interaction mechanism [1]. These updates strategically integrate the recommend actions generated by semantic reasoning as the policy references for joint decision, thereby generating the execution action policy. This ensures the dynamic responsiveness to evolving environmental states, achieving stable and operationally efficient control behaviors. The DRL framework is described in detail below.
  • Markov Decision Process (MDP) Formulation: The real-time action policy generation problem is formulated as an MDP, including the state space, action space, and reward. The state $\mathbf{s}$ consists of the collected multimodal data, including the vehicle’s own status, e.g., the velocity and textual modalities, and the environmental information, e.g., the visual and radar modalities. The action $\mathbf{a}$ is sampled from the advantage estimates over all feasible actions within the discrete action space, involving the driving maneuvers, e.g., turning left, turning right, proceeding straight, accelerating, decelerating, and maintaining constant speed. The reward is designed as the sum of all reward terms at time-step $t$, i.e., $R\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)=R_{t}^{\textrm{sfty}}+R_{t}^{\textrm{de}}+R_{t}^{\textrm{comf}}$, including the safety, driving efficiency, and comfort terms. Safety term penalizes the lane-change behavior that jeopardizes driving safety, with a penalty of $\delta_{1}$} when collision occurs. Driving efficiency term, including speed and lane-change rewards. Vehicle’s speed undergoes the linear normalization within a predefined target speed range. Specifically, the speed below the minimum threshold yields no reward, while the speed approaching the maximum value receives higher reward. When executing a lane-change maneuver under the environment safety constraint, EI vehicle receives a reward of $\delta_{2}$}. This reward encourages proactive and flexible driving policies, when slow moving vehicles are detected ahead. Comfort term provides a reward of $\delta_{3}$}, when staying close to the rightmost lane, promoting a preference for the rightmost lane and avoiding unnecessary lane changes. These terms enable the EI vehicle to learn an optimal policy $\pi^{*}$ that maximizes the expected return.
  • Environment Interaction and Training Process: At each time-step $t$, EI vehicle interacts with the environment and selects an action $\mathbf{a}_{t}$ from current state $\mathbf{s}_{t}$ based on a noisy network. Upon taking action $\mathbf{a}_{t}$, EI vehicle receives a reward $R\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)$ and the environment transits to next state $\mathbf{s}_{t+1}$. The transition is stored into the experience replay buffer, and current state $\mathbf{s}_{t}$ and next state $\mathbf{s}_{t+1}$ are fed into the evaluation network and target network, respectively, to obtain the Q-values for different actions. The action $\mathbf{a}_{t+1}$ corresponding to the maximum Q-value is selected from the evaluation network, and its Q-value is obtained from the target network. This value is then used to guide the training and update the evaluation network. The target network gradually adjusts its parameters through soft updates to reduce the instability during training.
Stage 3. Fusion Pipeline. With the first two stages, the final execution action is obtained through a policy pipeline, while the semantics output from the semantic pipeline is referenced during the generation process. However, due to the differences in decision-making of semantic and policy pipelines, their generated results may not exhibit the complete consistency. Without effectively identifying such differences, LMMs may encounter difficulties in maintaining alignment with the actual policy during long-term reasoning. To address this issue, the fusion pipeline is introduced to enable the action consistency discrimination and facilitate the backward updating of the policy, enabling continuous learning.
  • Step 1. Action Comparison: The action comparison step compares the action executed by the policy pipeline with the reasoning output of the semantic pipeline under the same states. If the two outcomes are consistent, this step indicates that the semantic reasoning has successfully aligned with the policy preference, requiring no further adjustment. Otherwise, the discrepancy is identified and constructed into a feedback sample.
  • Step 2. Feedback & Adaptation: When action inconsistency is detected, the step of feedback and adaptation sends the corresponding state-action pair as a feedback sample back to the semantic pipeline. The sample helps fine-tune the reasoning parameters and produce a new output. In the short term, it helps ensure real-time response and operational stability during execution. Over the long term, this enforces deep coupling between the semantics and policy pipelines, which enhances the generalization ability of EI driving system in complex environment.
III-B Benefits
The proposed framework introduces the following potential benefits for the EI driving system.
Enhanced Environmental Perception. The EI driving system uses multimodal heterogeneous data to predict road participants’ behavioral intention and its ultimate environmental understanding. This cross-modal information processing capability of EI vehicle transcends the static limitation of conventional perception, resulting in much better awareness of the environment.
Improved Decision-Making Reliability. The EI driving system dynamically adjusts the driving policy in real-time within dynamic environment, while rapidly identifying anomalous objects and executing strategic planning. This ensures robust state estimation and policy selection in complex, open-world scenarios, thereby enhancing the reliability of decision-making.
Strengthened Continual Learning Capability. During the training process, the EI driving system converts the action discrepancies into the learning signals, and fine-tunes them in combination with historical interactive data, enabling the policy adaptation to unknown scenarios during operation. The stronger environmental adaptability and improved sustainable learning capacity can be obtained.
III-C Use Cases
Within the proposed framework, several promising use cases are illustrated below.
Lane Change. As a challenging driving maneuver, lane-change requires EI vehicle to maintain stable move within its original lane, while simultaneously selecting and entering a target lane. Our framework is capable of reasoning about the lane-changing feasibility and generating the specific lane-change planning actions. This thereby enables the safe and efficient lane-change execution within complex traffic flows.
Ramp Merging. In the ramp merging scenarios, EI vehicle must safely and smoothly merge into the main traffic flow within constrained spatial and temporal conditions. Our framework enables the accurate traffic scene interpretation and the merging opportunity identification. Real-time merging decisions can also be generated to achieve seamless integration with dynamic traffic flows.
Intersection Negotiation. Intersection negotiation presents multiple interaction challenges for EI driving. Our framework can analyze traffic regulations, identify the intentions of other vehicles, e.g., conventional human-driven vehicles, and interpret pedestrian behaviors. Meanwhile, path planning and behavior coordination can be well performed within the road structure, thereby maintaining both traffic efficiency and safety in highly uncertain intersection environment.
Obstacle Avoidance. Obstacle avoidance involves the rapid detection and response to static and dynamic obstacles. Our framework can precisely identify the obstacle locations and motion states, and obtain immediate avoidance actions. Both the continuity of vehicle motion and operational safety can be thereby ensured under extreme scenarios.
IV A Case Study of Lane-Change Planning
This section presents a case study for lane-change planning to evaluate the performance of the proposed framework.
Scenario and Baselines. Consider a lane‑change scenario for an EI vehicle moving on a two‑lane road of $7\,\textrm{m}\times 3,000\,\textrm{m}$. In addition to this EI vehicle, $35$ conventional human-driven vehicles are uniformly distributed across this road. We use the fine‑tuned PaliGemma model [^1] https://github.com/hanker-zhu/DriveVLM-project., to provide the action recommendations. The PaliGemma model consists of a Transformer decoder and a Vision Transformer-based image encoder, with a total of 3 billion parameters. The text decoder is initialized with weights from the Gemma 2B model, and the image encoder uses a pre-trained SigLIP So400m Patch14 architecture. We compare the proposed framework with two baseline schemes: i) Integration of LMMs with the state-of-the-art DRL algorithms, including DDQN and DQN; ii) EI driving scheme without the assistance of LMMs (W/O LMM), where the semantic pipeline is not included, and the network structure of DRL approach remains identical as our framework.
Figure 5: Performance evaluation of the proposed dual-driven hybrid decision framework, compared with the baseline schemes including the state-of-the-art DRL algorithms and the particular case of W/O LMM: (a) Convergence performance of the adopted D3QN algorithm integrated with LMMs in the proposed framework; (b) Average reward versus the number of human-driven vehicles.
Dataset. Throughout the experiments, we employ the open-source nuScenes dataset [^2] https://www.nuscenes.org/nuscenes., which integrates the structured data from three modalities. These modalities include multi-view scene videos (local), bird’s-eye-view (BEV) map images (global), and multi-round Question Answering (QA) annotations. The nuScenes dataset consists of images from 1,000 driving scenes, captured across diverse urban and highway environments. Besides, it includes 18,000 reasoning-based QA pairs, and more than 50,000 reasoning steps. This ensures the model learns to generate structured rationales before arriving at a final answer. We utilize a subset of this dataset as the multimodal information input for LMMs. These data are used for the preliminary training of EI vehicle’s lane-change task.
Training and Hyperparameters. We adopt Low‑Rank Adaptation (LoRA) method for fine-tuning. The AdamW optimizer is utilized to reduce memory requirements and achieve rapid adaptation to driving semantics and actions. The fine-tuned model is integrated with the D3QN algorithm [1], for reducing erroneous interventions and outputting deterministic policy. The D3QN algorithm uses a four-layer fully connected neural network with two hidden layers, each containing 256 neurons and ReLU activation. The evaluation network updates its parameters via Adam optimizer. We set the reward-associated parameters as ${\delta_{1}}=-15$, ${\delta_{2}}=10$, and ${\delta_{3}}=2$, respectively. The experiment is configured with 3,000 training episodes, local batch size of 32, and learning rate of 0.001.
Results Evaluation. Fig. 5 presents the convergence performance of the adopted D3QN algorithm integrated with LMMs in our framework against the baselines. In DRL, rewards constitute the sole feedback signal during training and provide a direct measure of the policy quality. As observed, our framework exhibits superior convergence performance and ultimately achieves a higher average reward than all baselines. Notably, our framework outperforms both the DDQN and DQN algorithms by achieving an average reward gain of at least 19.47% and 31.07%, respectively. This superiority stems from the Dueling structure of D3QN, which decouples the state values and action advantages. In contrast, DDQN only addresses Q-value overestimation, while DQN’s single network structure is more prone to local optima, leading to limited exploration efficiency. We can also see that the W/O LMM scheme shows inferior convergence performance. This is because W/O LMM relies entirely on the algorithm for exploration. Due to the lack of the guidance from LMMs, W/O LMM fails to obtain the reward gains brought by action consistency, resulting in slower convergence and lower reward.
Fig. 5 illustrates the impact of number of conventional human-driven vehicles on the average rewards for our framework and the baselines. As shown, our framework consistently outperforms all baselines, demonstrating superior dynamic adaptability. Particularly, all schemes exhibit a trend of initial increase followed by a decrease. As the number of conventional vehicles increases, the average reward rises gradually, peaking at 35 vehicles, before declining. This is because the EI vehicle gains more lane-change opportunities due to the moderate increase in conventional vehicles, facilitating higher reward. However, an increasing number of conventional vehicles reduces driving maneuvering space, compelling the EI vehicle to adopt conservative driving behavior to avoid collision, thereby limiting reward growth.
V Future Research Directions
V-A Virtual and Reality Collaborative Training
By exploiting the high‑fidelity physical simulation environment and the world models, EI vehicle undergoes large‑scale pre‑training and policy optimization in virtual space, thereby achieving the controlled coverage of long‑tail scenarios. Fine-tuning with real-world data enables efficient transfer from virtual domain to real domain. The efficiency of task execution increases, while the risk factor associated with physical training decreases. Therefore, the implementation of the virtual and real collaborative training paradigm into the development process of EI driving seems to be a reasonably key research direction.
V-B Endogenous Security Defense
With the popularization of V2X communications, EI driving system is increasingly exposed to multiple cyber threats. Existing security mechanisms often struggle to promptly respond to such emerging risks. Endogenous security is the embedding of adaptive and self-evolving security solutions in the core architecture of the system with a security consideration from the design stage itself. This approach establishes the proactive defense paradigm, enabling the system to autonomously identify, resist, and mitigate various security threats. Therefore, integrating such security‑endogenous paradigm into the system architecture represent a promising research direction.
V-C Artificial General Intelligence (AGI)-Enabled Multi-Agent Collaboration
In open-world traffic scenarios, EI vehicle must perform real-time intent inference and engage in strategic interaction with other road participants. However, it is difficult for existing methods to simultaneously achieve interpretability, adaptability, and safety within the dynamic multi-agent interactions. Therefore, it is necessary to build an interpretable and adaptive multi-agent interaction framework through AGI, for realizing collaborative decision‑making within safety constraints. This represents a crucial research direction for advancing the large‑scale implementation of EI driving system.
VI Conclusion
In this article, we have proposed a semantics and policy dual-driven hybrid decision framework to achieve continuous learning and joint decision-making for enhancing EI driving. The proposed framework has merged LMMs, for semantic understanding and cognitive representation, and DRL, for real-time policy optimization. We also have provided the emerging opportunities this framework empowers, from the potential benefits to the transformative use cases. Furthermore, a case study, which aims for lane-change planning in an EI vehicle and conventional vehicle mixed scenario, has demonstrated the superiority of our framework in policy quality and dynamic adaptability. Concluding the article, some prospective research directions pertaining to the EI driving have been outlined for future exploration.

### v3

@@
 The framework merges LMMs for semantic understanding and cognitive representation, and deep reinforcement learning (DRL) for real-time policy optimization.
-We starts by introducing the foundational principles of EI driving and LMMs.
+We start by introducing the foundational principles of EI driving and LMMs.
 Moreover, we examine the emerging opportunities this framework enables, encompassing potential benefits and representative use cases.
@@
 Unlike conventional AI models, embodied AI emphasizes the interactions between embodied agents and their physical environment [8].
-By embedding embodied AI, self-driving vehicle can adapt to dynamic environmental changes and make intelligent decisions based on the interactive data.
-Such adaptability enables the vehicles to continuously optimize and evolve their driving policies through interaction in complex environment, thereby realizing the embodied intelligent (EI) driving paradigm.
-Currently, embodied AI is in the exploratory stage, used for real-time decision-making and action planning in dynamic and uncertain environment.
+By embedding embodied AI, self-driving vehicles can adapt to dynamic environmental changes and make intelligent decisions based on the interactive data.
+Such adaptability enables the vehicles to continuously optimize and evolve their driving policies through interaction in complex environments, thereby realizing the embodied intelligent (EI) driving paradigm.
+Currently, embodied AI is in the exploratory stage, used for real-time decision-making and action planning in dynamic and uncertain environments.
 Great potential has been particularly demonstrated in the transportation-related scenarios, including vehicular networks and rail transportation systems [9, 10, 11, 12].
@@
 Although these achievements are of great significance, as shown in Table I, a closed-loop design mechanism capable of continuous learning in response to environmental changes within dynamic traffic conditions has yet to be established.
-While LMMs enhance understanding of multimodal driving data, deep reinforcement learning (DRL) is required to facilitate joint decision-making in dynamic environment.
-The decision process often necessitates continuous interaction between EI vehicle and its environment.
+While LMMs enhance understanding of multimodal driving data, deep reinforcement learning (DRL) is required to facilitate joint decision-making in dynamic environments.
+The decision process often necessitates continuous interaction between EI vehicles and their environment.
 Therefore, incorporating the continuous learning and joint decision in dynamic scenarios into EI driving remains an emergent research frontier.
@@
 LMMs inherit and extend the capabilities of LLMs, traditionally designed for textual data processing.
-Through architectural adaptation and multimodal collaborative training via LMMs, EI vehicle effectively handles the collected multimodal data [13].
-These data processed and integrated into internal representation thereby enhance its generalization ability to new environment and tasks.
+Through architectural adaptation and multimodal collaborative training via LMMs, EI vehicles effectively handle the collected multimodal data [13].
+These data processed and integrated into internal representation thereby enhance its generalization ability to new environments and tasks.
 In contrast, DRL behaviors can be adjusted by real-time feedback, which results in adaptive policy development in dynamic operations [14, 15].
@@
 Such approach shows great promise in three illustrative application scenarios: i) Urban robotaxi, which is on-demand and driverless transport, faces tremendous challenges in negotiating highly dynamic traffic scenarios.
-Examples would be predicting a pedestrians intent, and avoiding a sudden obstacle in a narrow corridor.
+Examples would be predicting a pedestrian’s intent, and avoiding a sudden obstacle in a narrow corridor.
 ii) Autonomous freight operates in trunk-line logistics, ports, and industrial parks, where good performance is needed in complex geometrical road situations.
@@
 II An Overview of LMMs and EI Driving
-This section begins with the fundamental concept of EI driving, followed by the general framework of LMMs. Then, the necessities of integrating EI driving and LMMs is discussed.
+This section begins with the fundamental concept of EI driving, followed by the general framework of LMMs. Then, the necessities of integrating EI driving and LMMs are discussed.
 II-A Fundamentals of EI Driving
@@
 Multimodal instruction tuning aims to comprehend and carry out cross-modal instructions, which refers to building the pairs of instruction semantics to the output data.
-LMMs applies the In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning capability in a multimodal environment, enhancing their cross-modal generalization performance.
+LMMs apply the In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning capability in a multimodal environment, enhancing their cross-modal generalization performance.
 In complex and varied cross-modal tasks, performance enhancement is especially observed.
@@
 Within this DRL framework, the policy network parameters are continuously optimized through online interaction mechanism [1].
-These updates strategically integrate the recommend actions generated by semantic reasoning as the policy references for joint decision, thereby generating the execution action policy.
+These updates strategically integrate the recommended actions generated by semantic reasoning as the policy references for joint decision, thereby generating the execution action policy.
 This ensures the dynamic responsiveness to evolving environmental states, achieving stable and operationally efficient control behaviors.
@@
 The nuScenes dataset consists of images from 1,000 driving scenes, captured across diverse urban and highway environments.
-Besides, it includes 18,000 reasoning-based QA pairs, and more than 50,000 reasoning steps.
+Besides, it includes 18,000 reasoning-based Question Answering pairs, and more than 50,000 reasoning steps.
 This ensures the model learns to generate structured rationales before arriving at a final answer.
@@
 This approach establishes the proactive defense paradigm, enabling the system to autonomously identify, resist, and mitigate various security threats.
-Therefore, integrating such security‑endogenous paradigm into the system architecture represent a promising research direction.
+Therefore, integrating such security‑endogenous paradigm into the system architecture represents a promising research direction.
 V-C Artificial General Intelligence (AGI)-Enabled Multi-Agent Collaboration